simplify complex fraction: numerator is x/(x+2) and denominator is 1/x + 1/x+2. I think answer is x^2/2x+2 but not sure how to get it.
Answer by rothauserc(4718)

we are asked to
simplify complex fraction: numerator is x/(x+2) and denominator is 1/x + 1/x+2
note that 1/x +1/(x+2) = x+2+x/x(x+2) then we get
x/(x+2) * x(x+2)/2x+2
cancel the x+2 term and we have
x * x / 2(x+1) = x^2 / 2(x+1)